his grandpa's 84 and his mom is 46 how many years ago was grandpa twice her age.
Answer by Theo(13342)

let x = the number of years ago.

you get 84 - x = 2 * (46 - x)

that results in 84 - x = 92 - 2x

add 2x to both sides of this equation and subtract 84 from both sides of this equation to get:

x = 8

he's 84 now.
8 years ago he was 76

she's 46 now.
8 years ago she was 38

2 * 38 = 76

looks like we did good.

answer is the age of grandpa was 2 times the age of mom 8 years ago.